1. find the sum, if possible:

     1+3+1/3+1/9+1/27+.........


1
Expert's answer
2021-06-08T10:18:44-0400

Geometric series


"\displaystyle\sum_{i=0}^\infin ar^n=\dfrac{a}{1-r}, |r|<1"

We have "a=1, r=\dfrac{1}{3}, |r|=\dfrac{1}{3}<1." Then


"\displaystyle\sum_{i=0}^\infin (\dfrac{1}{3})^n=1+\dfrac{1}{3}+\dfrac{1}{9}+\dfrac{1}{27}+...=\dfrac{1}{1-\dfrac{1}{3}}=\dfrac{3}{2}"

Therefore


"1+3+\dfrac{1}{3}+\dfrac{1}{9}+\dfrac{1}{27}+...=3+\dfrac{3}{2}="

"=\dfrac{9}{2}=4.5"